OpenStudy (anonymous):

I'm stuck on this linear inequality. it says: A furniture company makes 2 table legs; one plain and one fancy. The plain legs take 2 hrs on a lathe and 1 hour sanding to make. The Fancy legs take 1 hour on a lathe and 4 hrs sanding to make. the warehouse has 4 lathes and 6 sanding machines, which are available 12 hrs per day. Each plain leg nets $3 in profit and each fancy legs nets $5. how can the company maximize profit? I was told this could be done with calculus, but how can it be done with algebra?

OpenStudy (anonymous):

How is this done in calculus when the question is so long winded?

OpenStudy (anonymous):

Dunno, I don't know calculus. This was a question in a college algebra book, which is what I'm trying to learn.

OpenStudy (anonymous):

This question should be done in algebra

OpenStudy (anonymous):

The Chapter is about linear inequalities and linear programming

OpenStudy (anonymous):

It would definitely help me if someone could answer it using algebraic principles since I don't know calculus

OpenStudy (anonymous):

The terms to list in linear programming P=Plain, F = Fancy, H = hours P= 1S+2L of (manhours) = total is 3h F= 4S+1L of (manhourse) = total is 5h Available resources: 4L+6S = 12h Plain profit = $3/leg Fancy profit = $5/leg

OpenStudy (anonymous):

Using algebra we can solve for max profit= Try to equate the available resources to P & F's profit. At same time include their production per leg per manhour